OpenStudy (anonymous):

What is the 10th term in the geometric sequence with t1 = 8 and r = 1.5? Show your steps and round to the nearest hundredth.

OpenStudy (anonymous):

How would you find the second and third terms?

OpenStudy (anonymous):

Do you know or are you asking me?

OpenStudy (anonymous):

\[T(n)=a(r)^n-1\]

OpenStudy (anonymous):

That's how you do the problem right? 8 * (1.5)^(10-1) = 8 * 1.5^(9). Just that the -1 in the thing you typed is up with the n.

OpenStudy (anonymous):

Yea